# Adapting Noise to Data: Generative Flows from 1D Processes

### v3

@@ Abstract
We introduce a general framework for constructing generative models using one-dimensional noising processes. Beyond diffusion processes, we outline examples that demonstrate the flexibility of our approach. Motivated by this, we propose a novel framework in which the 1D processes themselves are learnable, achieved by parameterizing the noise distribution through quantile functions that adapt to the data. Our construction integrates seamlessly with standard objectives, including Flow Matching and consistency models. Learning quantile-based noise naturally captures heavy tails and compact supports when present. Numerical experiments highlight both the flexibility and the effectiveness of our method.
@@ 1 Introduction
Flow-based generative models, especially score-based diffusion [pmlr-v37-sohl-dickstein15, SE2019], flow matching (FM) [albergo2023stochastic, lipman2023flow, liu2022rectified] and consistency models like the recently introduced inductive moment matching (IMM) [IMM2025], achieve state-of-the-art results in many applications. All these methods construct a probability flow from a simple latent distribution (noise) to a complex target (data) with a neural network trained to approximate this flow from limited target samples. In diffusion models, the score function directs a reverse-time SDE, while in FM, the velocity field is learned to compute trajectories via a flow ODE. Finally, consistency models like IMM learn to predict the jumps from noise to the data while factoring in the consistency of the flow trajectories. Usually, a Gaussian is used as latent distribution which causes difficulties when learning certain multimodal and heavy-tailed targets [HN2021, SBDD2022], see Figure 5 for a heavy-tailed example. There exist only few approaches to learn the noising process, [bartosh2025neuralflowdiffusionmodels] fit the forward diffusion process via a learned invertible map that is trained end-to-end, [kapusniak2024metric] use metric flow matching, i.e., a neural network to adapt the path to a underlying Riemannian metric. In a related approach, [sahoo2024mulan] learns a componentwise Gaussian noise schedule, input-conditioned so each component receives its own noise level. In the setting of sampling from unnormalized target densities, [Blessing2025UDB] learn the latent noise by optimizing the mean and covariance of a Gaussian prior, while [blessing2025endtoendlearninggaussianmixture] learn a Gaussian mixture prior, both are trained end to end. On the other hand [pandey2024heavytaileddiffusionmodels, zhang2024longtailed] design heavy-tailed diffusions using Student-$t$ latent distributions, and [HTDL2025] extend the framework to the family of $\alpha$-stable distributions.
@@
Figure 1: FM via optimal coupling with Gaussian noise (left) and our learned noise (right). Samples from the latent distribution are black and from generated ones are red. The left image shows their transportation paths in green. Starting from the learned latent drastically shortened the paths. (Zoom in for the paths in the bottom image).
In this paper, we present a new approach to adapt the latent distribution to the data by learning from its samples. The basic idea comes from the fact that all the above methods implicitly emerge as componentwise models. For example, denoting the target random variable by $\mathbf{X}_{0}$ and the latent by $\mathbf{X}_{1}\sim\mathcal{N}(0,I_{d})$, FM utilizes the process $\mathbf{X}_{t}=(X^{1}_{t},\ldots,X_{t}^{d})$ with the components $X^{i}_{t}=(1-t)X_{0}^{i}+tX_{1}^{i}$ employing one-dimensional Gaussians $X_{1}^{i}\sim\mathcal{N}(0,1)$. This motivated us to generally construct generative models from 1D processes and their quantile functions.
Given any appropriate 1D process we demonstrate how to learn the componentwise neural flow by the associated conditional velocity field. We give examples besides diffusion demonstrating the flexibility of our machinery, namely the Kac process arising from the 1D damped wave equation, see [DCFS2025, distillkac2025], and a process reflecting the Wasserstein gradient flow of the maximum mean discrepancy with negative distance kernel towards the uniform distribution. In contrast to diffusion, assuming a compactly supported target, these processes also have a compact support, leading to a better regularity of the corresponding velocity field. This inspired us to further adapt the process to the data and to learn the 1D noising process rather than choosing it manually. To this end, we exploit that 1D probability measures can be equivalently described by their quantile functions $Q^{i}:(0,1)\to\mathbb{R}$ which are monotone functions, and consider quantile processes $X_{t}^{i}=(1-t)X_{0}^{i}+tQ^{i}(U^{i})$, $i=1,\ldots,d$ with i.i.d. $U^{i}\sim\mathcal{U}[0,1]$ for $t\in[0,1]$. We learn the individual quantile functions $Q^{i}_{\phi}$, $i=1,\ldots,d$ such that their componentwise concatenation $\mathbf{Q}_{\phi}(\mathbf{U})\coloneqq(Q_{\phi}^{i}(U^{i}))_{i=1}^{d}$ is "close" to the data.
This inspired us to minimize
$W_{2}^{2}(\mu_{0},{\rm Law}(\mathbf{Q}_{\phi}(\mathbf{U}))),\quad\mu_{0}=\text{Law}(\mathbf{X}_{0}).$
with the Wasserstein distance $W_{2}$. We combine the learning of the latent $\mathbf{Q}_{\phi}(\mathbf{U})$ with the learning of the velocity field via optimal coupling FM. This allows us to effectively exploit the learned noise and drastically shorten the transport paths, as illustrated in Figure 1.
The simplicity of quantile functions give us a flexible tool, which enables us to simultaneously learn the noising process and apply the FM framework. Our quantile perspective can further be extended to fit into consistency models.
@@
1. We introduce a general construction method for generative neural flows by decomposing multi-dimensional flows into one-dimensional components. Ultimately, this allows us to work with one-dimensional noising processes in the FM framework.
2. We highlight three interesting noising processes for our framework: the Wiener process, the 1D Kac process and the 1D MMD gradient flow with negative distance kernel and uniform reference measure.
3. Based on the decomposition viewpoint, we propose to describe our 1D noising processes by their quantile functions. Via quantile interpolants, our framework can also be incorporated into consistency models.
4. Exploiting the simplicity of quantile functions, we propose to learn the quantile functions of the 1D noise simultaneously within the FM framework, aiming to fit the noise to the data. Numerical experiments demonstrate the high flexibility of our data-adapted latent noise and a shortening of transport paths.
Outline of the paper
In Section 2, we start with essentials for our method, namely absolutely continuous curves in Wasserstein spaces in Subsection 2.1 followed by conditional FM in Subsection 2.2. Then, in Subsection 2.3, we introduce quite general so-called mean-reverting processes which cover standard interpolation processes in FM, but will be required in their generality when dealing with the Kac flows and quantile interpolants for consistency models.
In Section 3, we propose to decompose the noising processes into one dimensional ones. Having concrete one-dimensional processes in mind as the Kac process or the Wasserstein gradient descent flow of a special mean discrepancy functional leading to a certain Uniform process, we provide a general construction method for building accessible conditional velocity fields in FM. Finally, we consider just a scaled process, where the Wiener process and the Uniform one are special cases of. This leads us immediately to the question which one-dimensional processes or more precisely which latent noise distribution should be chosen.
Section 4, which is a main contribution of our paper, aims to learn meaningful latent distributions from the data by learning the quantile functions belonging to one-dimensional noise.
In Subsection 4.1, we recall the relation between measures on $\mathbb{R}$ and their quantile functions, and construct quantile interpolants that may enable the integration of our framework with consistency models like inductive moment matching. Based on this, we explain in Subsection 4.2 how a data-adapted latent distribution can be learned jointly with the velocity field.
Numerical experiments demonstrating the advantages and the high potential of our approach are outlined for Section 5. On synthetic datasets, we first analyze the behavior of flow matching combined with the learned latent, and subsequently characterize its advantages and limitations on image datasets.
2 Flow Matching and Stochastic Processes
In this section, we first review absolutely continuous curves in Wasserstein spaces as basis of the subsequent FM method. Then we highlight quite general stochastic processes $(\boldsymbol{X}_{t})_{t}$ “interpolating” between a noising process $(\boldsymbol{Y}_{t})_{t}$ that starts in $\boldsymbol{Y}_{0}=0$ and ends in $\boldsymbol{Y}_{1}$ (our latent noise) and our target $\boldsymbol{X}_{0}$. In particular, we provide the relation between the corresponding vector fields.
2.1 Absolutely Continuous Curves in Wasserstein Space
@@
We start with a brief introduction of curves in Wasserstein spaces and basic ideas on flow matching. For more details we refer to [BookAmGiSa05] and [WS2025]. Let $(\mathcal{P}_{2}(\mathbb{R}^{d}),W_{2})$ denote the complete metric space of probability measures with finite second moments equipped with the Wasserstein distance
$W_{2}^{2}(\mu,\nu):=\min_{\pi\in\Pi(\mu,\nu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|x-y\|^{2}\mathop{}\!\mathrm{d}\pi(x,y)\big$.
@@
Here $\Pi(\mu,\nu)$ denotes the set of all probability measures on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ having marginals $\mu$ and $\nu$. The push-forward measure of $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ by a measurable map $\mathcal{T}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is defined by $\mathcal{T}_{\sharp}\mu:=\mu\circ\mathcal{T}^{-1}$. Let $I$ be an interval in $\mathbb{R}$, in this paper mainly $I=[0,1]$. A narrowly continuous curve $\mu_{t}:I\to\mathcal{P}_{2}(\mathbb{R}^{d})$ is absolutely continuous, iff there exists a Borel measurable vector field $v:I\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ with $\|v_{t}\|_{L_{2}(\mathbb{R}^{d},\mu_{t})}\in L_{2}(I)$ such that $(\mu_{t},v_{t})$ satisfies the continuity equation
@@
in the sense of distributions. If in addition $\int_{I}~\,\sup_{x\in B}\|v_{t}(x)\|+\mathrm{Lip}(v_{t},B)\mathop{}\!\mathrm{d}t<\infty$ for all compact $B\subset\mathbb{R}^{d}$, then the ODE
@@
Starting in the target distribution $\mu_{0}$ and ending in a simple latent distribution $\mu_{1}$, as usual in diffusion models, we can reverse the flow from the latent to the target distribution using just the opposite velocity field $v_{1-t}$ in the ODE (2). Thus, if somebody provides us with the velocity field $v_{t}$, we can sample from a target distribution by starting in a sample from the latent one and then applying our favorite ODE solver.
2.2 Flow Matching
If we do not have a velocity field donor, we can try to approximate (learn) the velocity field by a neural network $v_{t}^{\theta}$. Clearly, a desirable loss function would be
$\mathcal{L}(\theta)\coloneqq\mathbb{E}_{t\sim\mathcal{U}(0,1),\,x\sim\mu_{t}}\left[\left\|v_{t}^{\theta}(x)-v_{t}(x)\right\|^{2}\right].$ (3)
Unfortunately this loss function is not helpful, since we do not know the exact velocity field $v_{t}$ nor can sample from $\mu_{t}$ in the empirical expectation. However, employing the law of total probabilities, as done, e.g. in [lipman2023flow], we see that $\mathcal{L}(\theta)=\mathcal{L}_{\mathrm{CFM}}(\theta)+\text{const}$ with a constant not depending on $\theta$ and the Conditional Flow Matching (CFM) loss
$\mathcal{L}_{\mathrm{CFM}}(\theta)\coloneqq\mathbb{E}_{x_{0}\sim\mu_{0},\,t\sim\mathcal{U}(0,1),\,x\sim\mu_{t}(\cdot|x_{0})}\left[\left\|v_{t}^{\theta}(x)-v_{t}(x|x_{0})\right\|^{2}\right].$ (4)
The key difference is the use of the conditional flow $v_{t}(x|x_{0})$ with respect to a fixed sample $x_{0}$ from our target distribution. To summarize, all you need is a conditional flow model with accessible velocity field $v_{t}(x|x_{0})$ (at least along the flows trajectory), where you can easily sample from. Then you can indeed learn the velocity field $v_{t}$ of the general (non-conditional) flow and finally sample from the target by the reverse ODE (2).
2.3 Stochastic Processes and Velocity Fields
Consider a continuously differentiable (noising) process $(\mathbf{Y}_{t})_{t}$ with $\mathbf{Y}_{0}\equiv 0\in\mathbb{R}^{d}$ with associated velocity field $v_{t}=v^{\mathbf{Y}}_{t}(\cdot\mid 0)$ such that the pair $(\mu^{\mathbf{Y}}_{t},v^{\mathbf{Y}}_{t})$ satisfy the continuity equation (1), where $\mu^{\mathbf{Y}}_{t}$ is the law of $(\mathbf{Y}_{t})_{t}$ [^2] Existence of the velocity is given under weak assumptions by [WS2025] Theorem 6.3.. To construct a generative model we need to create a process $(\mathbf{X}_{t})_{t}$ which can start in any sample $x_{0}$ from the target measure $\mu_{0}$. Let $\mathbf{X_{0}}\sim\mu_{0}$. Following the lines in [DCFS2025], we define the mean‐reverting process by
$\mathbf{X}_{t}\;\coloneqq\;f(t)\,\mathbf{X}_{0}\;+\;\mathbf{Y}_{g(t)},\quad t\in[0,1],$ (5)
with smooth scheduling functions $f,g$
@@ 2.3 Stochastic Processes and Velocity Fields
Then we have $\mathbf{X}_{1}=\mathbf{Y}_{1}$, and by abuse of notation, the process $\mathbf{X}_{t}$ starts in $\mathbf{X}_{0}=\mathbf{X}_{0}$. Differentiation of (5) results in
$\dot{\mathbf{X}}_{t}=\dot{f}(t)\,\mathbf{X}_{0}\;+\;\dot{g}(t)\,\dot{\mathbf{Y}}_{g(t)}.$
The conditional velocity field of $\mathbf{X}_{t}$ is given by (see [WS2025, liu2022rectified])
@@
$=\mathbb{E}\bigl[\dot{f}(t)\,x_{0}+\dot{g}(t)\,\dot{\mathbf{Y}}_{g(t)}\,\bigm|\,\mathbf{Y}_{g(t)}=x-f(t)x_{0}\bigr]$
$=\dot{f}(t)\,x_{0}\,+\,\dot{g}(t)\,v^{\mathbf{Y}}_{g(t)}\bigl(x-f(t)x_{0}\mid 0\bigr).$ (7)
Now, the conditional flow matching loss (4) can be minimized regarding $\mathbf{X}_{0}\sim\mu_{0}$ and $\mathbf{X}_{t}\sim\mu_{t}$. Note that given a sample $x\sim(\mathbf{X}_{t}\mid\mathbf{X}_{0}=x_{0})$, we have $v^{\mathbf{X}}_{t}(x\mid x_{0})=\dot{f}(t)\,x_{0}\,+\,\dot{g}(t)\,v^{\mathbf{Y}}_{g(t)}\bigl(\mathbf{Y}_{g(t)}\mid 0\bigr)$. In general, $v^{\mathbf{Y}}$ might not be tractable, and only given as an conditional expectation of the time derivative $\dot{\mathbf{Y}}$. Yet, through our componentwise construction below, we will obtain easier access to it via its 1D components.
Remark 1 (Relation to FM and diffusion).
@@
Choosing $f(t)\coloneqq\alpha_{t}$, $g(t)\coloneqq\sigma_{t}^{2}$ and the standard Brownian motion $\mathbf{Y}_{t}=\mathbf{W}_{t}$, it holds the equality in distribution

### v4

# Adapting Noise to Data: Generative Flows from Learned 1D Processes

@@ Abstract
The default Gaussian latent in flow-based generative models poses challenges when learning certain distributions such as heavy-tailed ones. We introduce a general framework for learning data-adaptive latent distributions using one-dimensional quantile functions, optimized via the Wasserstein distance between noise and data. The quantile-based parameterization naturally adapts to both heavy-tailed and compactly supported distributions and shortens transport paths. Numerical results confirm the method’s flexibility and effectiveness achieved with negligible computational overhead.
Machine Learning, Generative Modeling, Heavy Tails, Flow Matching, Diffusion, Optimal Transport, Wasserstein, ICML
@@ 1 Introduction
Flow-based generative models have become a dominant paradigm in modern generative modeling. In particular, score-based diffusion models (Sohl-Dickstein et al., 2015; Song and Ermon, 2019; Song et al., 2021), flow matching (FM) methods (Albergo et al., 2023; Lipman et al., 2023; Liu, 2022), and more recently few-step approaches such as consistency-style models (Song et al., 2023; Geng et al., 2025; Boffi et al., 2025b) achieve state-of-the-art performance across a wide range of domains, including image synthesis and molecular generation (Hoogeboom et al., 2022), as well as discrete modalities such as text (Austin et al., 2023).
In flow-based models, the default choice of noise distribution is Gaussian, which can cause difficulties when learning, for example, multimodal or heavy-tailed target distributions; we refer to the extensive literature (Wiese et al., 2019; Hagemann and Neumayer, 2021; Salmona et al., 2022; Pandey et al., 2024; Ghane et al., 2025; Tam and Dunson, 2025). As a typical example, see the Neal’s funnel in Figure 2.
To enable sampling from heavy-tailed target distributions, recent state-of-the-art approaches typically replace the standard Gaussian reference with a heavy-tailed noise distribution. However, these methods still require manually tuning the tail behavior to match the data. For instance, (Pandey et al., 2024) proposes a Student-$t$ noise distribution with a tunable degrees-of-freedom parameter $\nu$. Similarly, (Shariatian et al., 2025b) extends the SDE framework to heavy-tailed settings by driving the dynamics with $\alpha$-stable noise.
@@
Figure 1: FM via optimal coupling with Gaussian noise (left) and our learned noise (right). Latent samples are shown in black, generated in red and transportation paths in green. Starting from the learned latent drastically shortens the paths.
In this paper, we propose a novel lightweight approach to design a noise distribution tailored to the data by learning it directly from samples. Motivated by the componentwise independent structure of the Gaussian and the Brownian motion, we construct multidimensional flows from one-dimensional ones. This enables us to address the latter ones by their quantile functions. While quantile functions can represent any 1D probability distribution, their monotonicity makes them particularly simple to parameterize, for example by using rational quadratic splines (Gregory and Delbourgo, 1982). More precisely, for quantile functions $Q^{i}$, $i=1,\ldots,d$, we consider the quantile processes
$X_{t}^{i}=(1-t)\,X_{0}^{i}+t\,Q^{i}(U^{i}),\qquad t\in[0,1],$ (1)
with i.i.d. $U^{1},\ldots,U^{d}\sim\mathcal{U}[0,1]$. We learn individual quantile functions $Q^{i}_{\phi}$ such that their componentwise concatenation $\mathbf{Q}_{\phi}(\mathbf{U})\coloneqq(Q_{\phi}^{1}(U^{1}),...,Q_{\phi}^{d}(U^{d}))$ is ”close” to the data by minimizing the Wasserstein distance between the data distribution and the noise. Note that this is a constrained optimization problem, since the noise is learned as a product of 1D measures. The simplicity of quantile functions gives us a flexible tool, which enables us to simultaneously learn the noising process and apply the FM framework.
Our approach allows us to i) avoid the limitations of Gaussian base noise, ii) eliminate manual fine-tuning of the noise distribution, and iii) drastically shorten the resulting transport paths as illustrated in Figure 1. Further, it can be extended to fit into few-step models, e.g. inductive moment matching (IMM) (Zhou et al., 2025; Boffi et al., 2025a) using more general mean reverting processes.
Figure 2: Sampling of Neal’s funnel with different latent distributions. [^2] Note that we used the independent coupling for training of these models. We also used z-score normalization.  Left to Right: uniform on [1,1][-1,1], standard Gaussian, Student-T (with parameters $(20,4)$ inspired by the choice in (Pandey et al., 2024)) and our learned distribution. The last two heavy-tailed noises perform significantly better.
@@
1. We introduce a general construction method for generative neural flows by decomposing multi-dimensional flows into one-dimensional components and applying a general mean reverting process. Ultimately, this allows us to work with 1D noising processes in the FM framework.
2. We examine interesting 1D noising processes besides Brownian motion: the physics-inspired Kac process and the MMD gradient flow, leading to compactly supported noise and a better regularity of the FM velocity field.
3. We show how our framework can be generalized for few-step models via so-called quantile interpolants. Furthermore, we offer an alternative method to construct multi-dimensional processes from one-dimensional ones using radially symmetric processes.
4. As main contribution, we propose to learn the 1D noise distributions within the FM framework in a data adapted way, by parameterizing them through quantile functions and employing the Wasserstein distance.
5. Numerical experiments demonstrate that our method efficiently handles diverse marginal structures including heavy-tailed, compact, and multi-modal distributions. Learned quantiles shorten transport paths by capturing per-coordinate structure, while delegating cross-dimensional dependencies to the velocity field.
2 Background on Flow Matching
In general, flow models can be described mathematically via curves in Wasserstein space, see, e.g. (Ambrosio et al., 2008) and (Lipman et al., 2024; Wald and Steidl, 2025). In the following, we denote the target distribution by $\mu_{0}\in\mathcal{P}(\mathbb{R}^{d})$ and the noise distribution by $\mu_{1}\in\mathcal{P}(\mathbb{R}^{d})$.
2.1 Curves in Wasserstein Spaces
@@
Let $(\mathcal{P}_{2}(\mathbb{R}^{d}),W_{2})$ denote the complete metric space of probability measures with finite second moments equipped with the Wasserstein distance
$W_{2}^{2}(\mu,\nu):=\min_{\pi\in\Pi(\mu,\nu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|x-y\|^{2}\,\mathrm{d}\pi(x,y)\big$.
@@
Here $\Pi(\mu,\nu)$ denotes the set of all probability measures on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ having marginals $\mu$ and $\nu$. By $\pi_{o}\in\Pi(\mu,\nu)$ we denote the minimizer of the right-hand side. The push-forward measure of $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ by a measurable map $\mathcal{T}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is defined by $\mathcal{T}_{\sharp}\mu:=\mu\circ\mathcal{T}^{-1}$. A narrowly continuous curve $\mu_{t}:[0,1]\to\mathcal{P}_{2}(\mathbb{R}^{d})$ is absolutely continuous, iff there exists a Borel measurable vector field $v:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ with $\|v_{t}\|_{L_{2}(\mathbb{R}^{d},\mathbb{R}^{d},\mu_{t})}\in L_{2}([0,1])$ such that $(\mu_{t},v_{t})$ satisfies the continuity equation
@@
in the sense of distributions.
2.2 From the Continuity Equation to the ODE
If the vector field fulfills $\int_{0}^{1}~\,\sup_{x\in B}\|v_{t}(x)\|+\mathrm{Lip}(v_{t},B)\,\mathrm{d}t<\infty$ for all compact $B\subset\mathbb{R}^{d}$, then the ODE
@@ 2.2 From the Continuity Equation to the ODE
Starting in the target distribution $\mu_{0}$ and ending in a simple latent distribution $\mu_{1}$, as usual in diffusion models, we can reverse the flow from the latent to the target distribution in inference using just the opposite velocity field $v_{1-t}$ in the ODE (3). If the velocity field $v_{t}$ is learned, we can sample from the target distribution by starting in a sample from the latent one and then applying our favorite ODE solver in (3).
2.3 Flow Matching: Learning the Velocity Field
Flow matching learns the velocity field $v_{t}$ with a neural network $v_{t}^{\theta}$ by minimizing the loss
$\mathcal{L}(\theta)\coloneqq\mathbb{E}_{t,x\sim\mu_{t}}\left[\left\|v_{t}^{\theta}(x)-v_{t}(x)\right\|^{2}\right]$ (4)
with uniformly sampled time in $[0,1]$. Indeed, this untractable loss can be rewritten as $\mathcal{L}(\theta)=\mathcal{L}_{\mathrm{CFM}}(\theta)+\text{const}$ with the conditional flow matching (CFM) loss
$\mathcal{L}_{\mathrm{CFM}}(\theta)\!\coloneqq\!\mathbb{E}_{t,x_{0}\sim\mu_{0},x\sim\mu_{t}(\cdot|x_{0})}\left[\left\|v_{t}^{\theta}(x)-v_{t}(x|x_{0})\right\|^{2}\right].$
Remark 2.1 (Flow Matching with Couplings.).
Consider a coupling $\pi\in\Pi(\mu_{0},\mu_{1})$ and the induced curve $\mu_{t}\coloneqq(e_{t})_{\sharp}\pi$ with $e_{t}(x,y):=(1-t)x+ty$. Then it is known that $\mathcal{L}_{\mathrm{CFM}}(\theta)$ can be rewritten as
$\mathcal{L}_{\mathrm{CFM}}(\theta)\!=\!\mathbb{E}_{t,\,(x,y)\sim\pi}\Big[\big\|v^{\theta}_{t}\big((1-t)x+ty\big)-(y-x)\big\|_{2}^{2}\Big].$ (5)
Taking the optimal coupling $\pi_{o}$ instead of the usually applied independent coupling $\pi=\mu_{0}\otimes\mu_{1}$, leads to a reduced variance in training and both shorter and straighter paths, see (Tong et al., 2024; Pooladian et al., 2023).
2.4 Velocity Fields via Stochastic Processes
Absolutely continuous curves and their velocity fields can be alternatively derived via stochastic processes: for a probability space $(\Omega,\Sigma,\mathbb{P})$ and a differentiable stochastic processes $(\mathbf{X}_{t})_{t\in[0,1]}$ with $\mathbf{X}_{t}\in L_{2}(\Omega,\mathbb{R}^{d},\mathbb{P})$, we have that $(\mu_{t},v_{t})$ with
$\mathbf{X}_{t}\sim\mu_{t}\coloneqq\mathbf{X}_{t,\sharp}\mathbb{P}\quad\text{and}\quad v_{t}\coloneqq\mathbb{E}[\mathbf{\dot{X}}_{t}|\mathbf{X}_{t}=\cdot]$ (6)
satisfies the continuity equation. We introduce a special mean reverting process which generalizes previous standard concepts, see Remark 2.2.
Mean Reverting Processes. Let us consider a continuously differentiable (noising) process $(\mathbf{N}_{t})_{t}$ with $\mathbf{N}_{0}\equiv 0\in\mathbb{R}^{d}$ and $\mathbf{N}_{t}\sim\mu^{\mathbf{N}}_{t}$ with associated velocity field $v^{\mathbf{N}}_{t}$. In the next section, we will show how to get access to such velocity fields. Then we define the mean‐reverting process by
$\mathbf{X}_{t}\;\coloneqq\;f(t)\,\mathbf{X}_{0}\;+\;\mathbf{N}_{g(t)},\quad t\in[0,1],$ (7)
with smooth scheduling functions $f,g$ fulfilling
@@ 2.4 Velocity Fields via Stochastic Processes
Then $\mathbf{X}_{1}=\mathbf{N}_{1}$ and the process $\mathbf{X}_{t}$ starts in $\mathbf{X}_{0}$. Differentiation of (7) results in
$\dot{\mathbf{X}}_{t}=\dot{f}(t)\,\mathbf{X}_{0}\;+\;\dot{g}(t)\,\dot{\mathbf{N}}_{g(t)}.$
Then, the conditional velocity field of $\mathbf{X}_{t}$ is given by
@@
$=~\mathbb{E}\bigl[\dot{f}(t)\,x_{0}+\dot{g}(t)\,\dot{\mathbf{N}}_{g(t)}\,\bigm|\,\mathbf{N}_{g(t)}=x-f(t)x_{0}\bigr]$
$=~\dot{f}(t)\,x_{0}\,+\,\dot{g}(t)\,v^{\mathbf{N}}_{g(t)}\bigl(x-f(t)x_{0}\bigr).$ (9)
Now, the conditional flow matching loss $\mathcal{L}_{\text{CFM}}$ can be minimized regarding $\mathbf{X}_{t}\sim\mu_{t}$: namely sampling $y\sim\mathbf{N}_{g(t)}$ we obtain a sample $x\sim(\mathbf{X}_{t}\mid\mathbf{X}_{0}=x_{0})$ from (7) satisfying
$v^{\mathbf{X}}_{t}(x\mid x_{0})=\dot{f}(t)\,x_{0}\,+\,\dot{g}(t)\,v^{\mathbf{N}}_{g(t)}\bigl(y\bigr).$
Remark 2.2 (Relation to FM and Diffusion).
@@
Choosing $f(t)\coloneqq\alpha_{t}$, $g(t)\coloneqq\sigma_{t}^{2}$ and the standard Brownian motion $\mathbf{N}_{t}=\mathbf{W}_{t}$, it holds the equality in distribution